OpenStudy (123ab456c):

Which ratio correctly compares 120 cm to 80 mm? A. 3 to 20 B. 3 to 2 C. 15 to 1 D. 150 to 1

jhonyy9 (jhonyy9):

your opinion ?

jhonyy9 (jhonyy9):

but first you need to convert 80 mm in cm so how many will be ?

jhonyy9 (jhonyy9):

80 mm = 8 cm so there 120/8 = 30/2 = 15/2 - hope you understand it

OpenStudy (anonymous):

Balance each unit to either CM or MM

OpenStudy (mathmale):

the original ratio is \[\frac{ 120.cm }{ 80.mm }.\] Multiply this by either the conversion formula for converting cm to mm or mm to cm. (Your answer should be the same in either case.)

OpenStudy (anonymous):

Then divide both sides to their smallest unit